# Learning to Insert for Constructive Neural Vehicle Routing Solver

### v2

@@ A.1 Constructive NCO Methods
Constructive NCO methods typically arrange the input nodes into high-quality solutions autoregressively. This paradigm is inspired by the sequence-to-sequence framework [23] in natural language processing (NLP), where a model encodes an input sequence into an embedding and then autoregressively generates an output sequence based on this representation. Vinyals et al. [6] pioneered this adaptation through their Pointer Networks (Ptr-Nets), which established the first sequence-to-sequence framework specifically for neural combinatorial optimization. In this paradigm, the neural model learns to predict the probability of appending the unvisited nodes to the partial solution until all nodes are visited. We call this paradigm Learning to Append (L2C-Append). Subsequent research has further advanced L2C-Append NCO methods, with improvements focusing on model training techniques [7], extension to solve other COPs Nazari et al. [8], and architectural enhancements [9]. Among them, the Attention Model [9] introduces the Transformer-based model structure [30], which achieves outstanding performances over various VRPs with up to 100 nodes. Various AM-based improved variants are proposed to narrow the performance gap to classical heuristic solvers [49, 10, 50, 51, 11, 52, 53, 54, 55, 13, 14, 56, 33, 57, 34, 58, 59, 17, 22, 60, 61]. Nonetheless, L2C-Append NCO methods are fundamentally limited by their restriction to appending new nodes, which can lead to suboptimal solutions due to early greedy choices. In contrast, the insertion-based paradigm offers greater flexibility for VRPs by sequentially inserting unvisited nodes into any valid position [24, 25], potentially yielding higher-quality solutions. Despite these benefits, the insertion-based constructive approach has remained unexplored in NCO for VRPs in the last decade. In this work, we introduce L2C-Insert, a novel learning to insert framework, to investigate the potential of this paradigm for constructive NCO.
@@ A.2 Non-constructive NCO Methods
This series of methods can be broadly categorized into three main groups: (1) Heatmap-based methods: These methods learn graph neural networks to predict a probability distribution, or heatmap, indicating the likelihood of each edge being part of the optimal solution for a given problem instance. Once the heatmap is obtained, it is used to guide subsequent decoding techniques such as greedy search, beam search [62], Monte Carlo Tree Search (MCTS) [37, 63, 64], 2-opt [44, 45], or gradient search [65] to generate the solution. While this approach can successfully solve TSP instances of various sizes, its application to other routing problems like CVRP is generally more challenging. (2) Improvement-based methods: These approaches iteratively refine solution quality. They employ neural models to predict regions within current solutions where heuristic operators, such as k-opt [48] and swap [66], can be applied to improve the solution quality [67, 68, 69, 70, 71, 32]. (3) Two-stage Method: Two-stage methods: These methods address the problem through an iterative divide-and-conquer process. In each iteration, the problem is first decomposed into multiple subproblems using rule-based techniques or neural models. Subsequently, neural solvers (e.g., AM [9]) or classical solvers (e.g., LKH3 [42] or HGS [36]) are used to solve these subproblems. The resulting partial solutions are then merged to form a complete solution [72, 70, 38, 27, 34]. Beyond these primary categories, other non-constructive NCO approaches include learning to augment classical solvers with neural components [73] or solving VRPs at a route level [74]. However, this paper primarily focuses on autoregressive constructive NCO methods that do not require heuristic operators or external solvers.
@@ B.2 Decoder
$\mathbf{e}_{i}$ =$=\begin{cases}[\mathbf{h}_{i},\mathbf{h}_{i+1}]&i\in\{1:l-1\},\\ [\mathbf{h}_{l},\mathbf{h}_{1}],&i=l,\end{cases}$ (5)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the horizontal concatenation operator. We note that both $\mathbf{h}_{1}$ and $\mathbf{h}_{l}$ are the depot embeddings. To incorporate the remaining capacity information, the decoder updates each position embedding by horizontally concatenating it with the remaining capacity of the route where the position is located, increasing the embedding dimension by 1. Subsequently, a linear projection $W_{2}\in\mathbb{R}^{{(2d+1)}\times d}$ is applied to the updated position embedding to project it to the embedding with the same dimension as the node embedding. The set of output position embeddings is denoted as $H_{e}=\{\mathbf{e}_{\pi_{i}}\},i\in\{1:t-1\}$.
Then, these embeddings together are refined by $L$ stacked attention layers. The first layer receives input $X^{0}=\{\mathbf{h}_{c},H_{e},H_{u}\}$, with each subsequent $j$-th attention layer processing the output from the $(j-1)$-th layer, culminating in the final output $X^{(L)}$. This process can be denoted as:
@@
Where the formulation of “AttnLayer” is detailed in Appendix G. Then, a linear projection and softmax function are applied to $X^{(L)}$ to compute the insertion probabilities of the current node into all valid positions. The masking operation assigns $\infty$- ∞ to invalid positions (those corresponding to the current node and non-visited nodes) to ensure proper probability normalization over valid insertion locations. The inserting probability $p_{i}$ is calculated as:
@@
$a_{i}$ =$=\begin{cases}x_{i}&\forall i\in\{1:l\}\text{ and the remaining capacity of the route where the position is located}\\ &\text{is not less than the current node's demand,}\\ -\infty&\text{otherwise,}\end{cases}$
@@ B.3 Implementation of Insertion-based Local Reconstruction on CVRP
  • Insertion-based Reconstruction The model inserts the unvisited nodes into the incomplete solution node by node. Initially, the last node is randomly selected from the incomplete solution. The subsequent procedures are the same as those for a greedy rollout. When all nodes are visited, a new complete solution is generated. We compare the previous solution with the new one, accepting the higher-quality solution for the next destruction-reconstruction cycle. This iterative process continues until the computational budget (e.g., maximum iterations or time) is reached.
@@ Appendix C Implementation Details of Baselines
For the heatmap-based methods, however, we encountered issues running their GPU-version code, such as incompatible issues when installing the required Pytorch extension library packages. Therefore, we opt to directly use their reported results. For SO-mixed, as their source codes are not released available, we choose to directly report their original results. However, we use the same test datasets for fair comparison (all these methods used the same TSP1K and TSP10K test datasets from Fu et al. [37]). For other methods, if not otherwise stated, we directly run their source codes with the default settings.
@@ Appendix D Hyperparameter Studies
We conduct hyperparameter studies for the values of neighborhood size $k$ and destruction size $\alpha$ by running L2C-Insert using local reconstruction with $I$=50 iterations.
@@ E.1 Effect of Distance Criterion for CVRP
We test two classic distance criteria, Euclidean Distance and Polar Angle. For two customer nodes $\mathbf{p}=(p_{1},p_{2})$ and $\mathbf{q}=(q_{1},q_{2})$, Euclidean Distance is defined as:$d_{E}(\mathbf{p},\mathbf{q})=\sqrt{(p_{1}-q_{1})^{2}+(p_{2}-q_{2})^{2}}$. For Polar Angle, we must introduce the depot node $\mathbf{d}_{c}=(d_{1},d_{2})$ to serve as the Polar center. Then it is defined as: $\theta_{P}(\mathbf{p},\mathbf{q})=\left|\theta_{\mathbf{p}}-\theta_{\mathbf{q}}\right|$, where $\theta_{\mathbf{p}}=\arctan\left(\frac{p_{2}-d_{2}}{p_{1}-d_{1}}\right)$ and $\theta_{\mathbf{q}}=\arctan\left(\frac{q_{2}-d_{2}}{q_{1}-d_{1}}\right)$ are the angles formed by the lines connecting the depot $\mathbf{d}_{c}$ to the customer nodes $\mathbf{p}$ and $\mathbf{q}$ with respect to the horizontal axis.
@@ E.5 Performance of L2C-Insert using Different Backbone Models
Appendix F Visualization

### v3

@@ A.1 Constructive NCO Methods
Constructive NCO methods typically arrange the input nodes into high-quality solutions autoregressively. This paradigm is inspired by the sequence-to-sequence framework [27] in natural language processing (NLP), where a model encodes an input sequence into an embedding and then autoregressively generates an output sequence based on this representation. Vinyals et al. [6] pioneer this adaptation through their Pointer Networks (Ptr-Nets), which establishes the first sequence-to-sequence framework specifically for neural combinatorial optimization. In this paradigm, the neural model learns to predict the probability of appending the unvisited nodes to the partial solution until all nodes are visited. We call this paradigm Learning to Append (L2C-Append). Subsequent research has further advanced L2C-Append NCO methods, with improvements focusing on model training techniques [7], extension to solve other COPs Nazari et al. [56], and architectural enhancements [19]. Among them, the Attention Model [19] introduces the Transformer-based model structure [36], which achieves outstanding performances over various VRPs with up to 100 nodes. Various AM-based improved variants are proposed to narrow the performance gap to classical heuristic solvers [57, 21, 58, 59, 60, 61, 62, 63, 64, 22, 23, 65, 39, 66, 40, 67, 68, 69, 26, 70, 71]. For example, PolyNet [33] enhances exploration by learning a diverse set of complementary solution strategies in a single model. By conditioning the solution generation process on a simple bit vector, PolyNet can generate diverse solutions without relying on handcrafted rules or forced initial moves. Nonetheless, L2C-Append NCO methods are fundamentally limited by their restriction to appending new nodes, which can lead to suboptimal solutions due to early greedy choices. In contrast, the insertion-based paradigm offers greater flexibility for VRPs by sequentially inserting unvisited nodes into any valid position [28, 29], potentially yielding higher-quality solutions. Despite these benefits, the insertion-based constructive approach has remained unexplored in NCO for VRPs in the last decade. In this work, we introduce L2C-Insert, a novel learning to insert framework, to investigate the potential of this paradigm for constructive NCO.
@@ A.2 Non-constructive NCO Methods
This series of methods can be broadly categorized into three main groups: (1) Heatmap-based methods: These methods learn graph neural networks to predict a probability distribution, or heatmap, indicating the likelihood of each edge being part of the optimal solution for a given problem instance. Once the heatmap is obtained, it is used to guide subsequent decoding techniques such as greedy search, beam search [72], Monte Carlo Tree Search (MCTS) [43, 73, 74], 2-opt [51, 52], or gradient search [75] to generate the solution. While this approach can successfully solve TSP instances of various sizes, its application to other routing problems like CVRP is generally more challenging. (2) Improvement-based methods: These approaches iteratively refine solution quality. They employ neural models to predict regions within current solutions where heuristic operators, such as k-opt [55] and swap [76], can be applied to improve the solution quality [77, 78, 79, 80, 81, 38]. (3) Two-stage Method: Two-stage methods: These methods address the problem through an iterative divide-and-conquer process. In each iteration, the problem is first decomposed into multiple subproblems using rule-based techniques or neural models. Subsequently, neural solvers (e.g., AM [19]) or classical solvers (e.g., LKH3 [49] or HGS [42]) are used to solve these subproblems. The resulting partial solutions are then merged to form a complete solution [82, 80, 44, 32, 40]. Beyond these primary categories, other non-constructive NCO approaches include learning to augment classical solvers with neural components [83] or solving VRPs at a route level [84]. A different approach is Neural Deconstruction Search (NDS) [85]. NDS uses a learned neural policy to iteratively deconstruct a solution by selecting customers to remove. These customers are then reinserted one by one using a rule-based greedy heuristic algorithm. In this paper, we develop a constructive neural model that learns the insertion-based heuristic, building solutions autoregressively.
@@ B.2 Decoder
[…]
where $\cdot$$\cdot$] is the horizontal concatenation operator. We note that both $\mathbf{h}_{1}$ and $\mathbf{h}_{l}$ are the depot embeddings. To incorporate the remaining capacity information, the decoder updates each position embedding by horizontally concatenating it with the remaining capacity of the route where the position is located, increasing the embedding dimension by 1. Subsequently, a linear projection $W_{2}\in\mathbb{R}^{{(2d+1)}\times d}$ is applied to the updated position embedding to project it to the embedding with the same dimension as the node embedding. The set of output position embeddings is denoted as $H_{e}=\{\mathbf{e}_{\pi_{i}}\},i\in\{1:t-1\}$.
Then, these embeddings together are refined by $L$ stacked attention layers. The first layer receives input $X^{0}=\{\mathbf{h}_{c},H_{e},H_{u}\}$, with each subsequent $j$-th attention layer processing the output from the (j1)(j-1)-th layer, culminating in the final output $X^{(L)}$. This process can be denoted as:
@@
Where the formulation of “AttnLayer” is detailed in Appendix N. Then, a linear projection and softmax function are applied to $X^{(L)}$ to compute the insertion probabilities of the current node into all valid positions. The masking operation assigns -$\infty$ to invalid positions (those corresponding to the current node and non-visited nodes) to ensure proper probability normalization over valid insertion locations. The inserting probability $p_{i}$ is calculated as:
@@
[…]
@@ B.3 Implementation of Insertion-based Local Reconstruction on CVRP
  • Insertion-based Reconstruction The model inserts the unvisited nodes to the incomplete solution node by node. Initially, the last node is randomly selected from the incomplete solution. The subsequent procedures are the same as those for a greedy rollout. When all nodes are visited, a new complete solution is generated. We compare the previous solution with the new one, accepting the higher-quality solution for the next destruction-reconstruction cycle. This iterative process continues until the computational budget (e.g., maximum iterations or time) is reached.
@@ Appendix C Implementation Details of Baselines
For the heatmap-based methods, we directly use their reported results. For SO-mixed, as their source codes are not released available, we choose to directly report their original results. However, we use the same test datasets for fair comparison (all these methods used the same TSP1K and TSP10K test datasets from Fu et al. [43]). For other methods, if not otherwise stated, we directly run their source codes with the default settings.
@@ Appendix D Hyperparameter Studies
We conduct hyperparameter studies for the values of neighborhood size $k$ and destruction size $\alpha$ by running L2C-Insert using local reconstruction with II=50 iterations.
@@ E.1 Effect of Distance Criterion for CVRP
[…]
@@ E.5 Performance of L2C-Insert using Different Backbone Models
Appendix F Discussion on the Advantage of the Insertion-based Construction Paradigm in Search
The iterative local reconstruction is a significant factor in achieving the final reported performance. We believe the key innovation of our work, the insertion-based policy, is what unlocks the true potential of this search procedure. The policy’s inherent flexibility is a key enabler for the search’s effectiveness.
1. The advantage of insertion-based search
As we discuss in Section 3.3, conventional appending-based models face a fundamental limitation when used with local search. When nodes are removed during the destruction phase, they can only be re-appended to the end of the sequence during reconstruction. This rigidity makes it difficult for the model to escape local optima. For instance, reconnecting two nodes that are geometrically close but far apart in the sequence—a common requirement for improving a tour—is highly improbable.
Our L2C-Insert framework directly overcomes this bottleneck. By allowing a node to be re-inserted at any valid position, our model facilitates a much more powerful and effective exploration of the solution space during the reconstruction phase. This flexibility is the core reason why the destroy-and-reconstruct loop is effective in our framework.
2. Direct Ablation Study: Isolating the Policy’s Contribution
To provide a direct and fair comparison that disentangles the policy’s contribution, we conduct the ablation study presented in Table 4 in the main paper. In this experiment, we compared our L2C-Insert model against an L2C-Append model (which uses a similar architecture but an appending policy) within the exact same local reconstruction framework.
The results show that our L2C-Insert with only 200 iterations significantly outperforms the L2C-Append model that uses 1000 iterations. This finding directly demonstrates that the superior performance stems from the intrinsic quality and flexibility of the insertion policy itself, and not just from the application of a search method.
Appendix G Clarification on Fair Comparison and Calculating Costs
1. Clarifying Iteration Settings of Baselines
To clarify, we summarize the iteration settings of iterative methods in the table below. From this table, we can observe that he number of iterations varies significantly. For instance, NeurOpt utilizes 10,000 lightweight k-opt iterations, while two-stage methods like UDC and SO-mixed perform 50-100 divide-and-conquer iterations.
Table 13: Iteration settings for baselilne methods. "Reviser-nn" refers to a solver designed to repair a local solution of length $n$.
Method Problem Number of iteration
NeurOpt TSP & CVRP 10000 (k-opt)
SO-mixed TSP1K 50 (Reviser-100/50)
TSP10K 100 (Reviser-100/50)
UDC TSP & CVRP 50 for Reviser-100 (with 50 initial solutions)
GLOP TSP 100/50/10 for Reviser-100/50/20
CVRP 5 for LKH3 as Reviser-20
2. A Note on Comparing Computational Budgets
We would like to respectfully clarify that directly comparing ’iteration counts’ may be challenging, as the definition and computational cost of an "iteration" differ substantially across methods:
  • Improvement-based methods like NeurOpt perform very fast, local search operations (e.g., k-opt) in each iteration.
  • Two-stage methods like GLOP and UDC involve multiple sub-problem solvers per iteration or perform the destroy-and-repair on multiple initial solutions, and also may call a powerful heuristic solver like LKH3 internally.
  • Our L2C-Insert performs local reconstruction only on a single solution using one neural model.
For this reason, the total inference time may offer a more holistic and equitable basis for comparing the computational budget. Our analysis is therefore centered on the solution quality (optimality gap) achieved within a given runtime. Our experimental results, when viewed in terms of total runtime, seem to support this perspective. For example, in Table 1 (TSP1K), our L2C-Insert (I=1000) achieves a 0.481% gap in 46.4 minutes. In contrast, a strong baseline like LEHD+RRC1000 requires 5.2 hours to achieve a 0.731% gap. This suggests that our method appears to offer a favorable trade-off between time and solution quality.
Appendix H Total Inference Time and Average Inference Time Per Instance
The apparent decrease in total inference time for larger instances is due to the different number of instances in the test datasets for each problem size.
Following the standard evaluation protocols established in prior NCO research [22, 43], the test sets for larger problems contain significantly fewer instances. For example, our TSP100 test set contains 10,000 instances, while the TSP10K and TSP100K sets contain only 16 instances each. The total inference times reported in Tables 1 and 2 of our main paper reflect the time required to solve the entire dataset.
To provide a clearer picture, the table below breaks down the total inference time into the average time per instance.
Table 14: Add caption
TSP100 TSP1K TSP10K TSP100K
Dataset Size 10000 instances 128 instances 16 instances 16 instances
L2C-Insert Total Inference Time 8.0h 46.4m 16.64m 26.07m
L2C-Insert Inference Time / instance 2.9s 21.8s 62.4s 97.8s
As the table clearly shows, the average inference time per instance for L2C-Insert scales with the problem size as expected, increasing from 2.9 seconds for TSP100 to 97.8 seconds for TSP100K.
Appendix I Clarification on Method Positioning and Technical Contribution
Regarding the positioning of our method (construction heuristic vs. improvement heuristic) and its technical contributions, we make the following discussion to clarify and strengthen the core value of our paper.
The "insertion-based construction" is a well-established concept in the traditional field of Operations Research. The core contribution of our work is not the invention of the ’insertion’ operation itself, but rather being the first learning-driven, insertion-based framework specifically designed for solution construction in the field of NCO for solving the vehicle routing problem, which is currently dominated by the ’appending-based’ paradigm.
The make the technical contribution of this paper clear, we elaborate from the following three perspectives:
1. The Essence of Our Core Method: A Pure Insertion-based Construction Method
Our core model, L2C-Insert, is fundamentally a construction-based model in its design. Its task is to learn a policy pp(position∣node, partial tour) to predict the best insertion position, given an incomplete solution (partial solution) and a node to be inserted. This process starts from an initial solution (e.g., one node) and builds a complete solution by sequentially adding unvisited nodes, which fully aligns with the definition of a construction-based neural model.
  • Model Architecture and Training: As described in Sections 3.1 and 3.2 of the paper, our model architecture (especially the decoder) and the supervised learning scheme are designed to learn this single-step constructive action. The model learns to master the optimal "insertion" decision at each step by reconstructing a given optimal solution. This is fundamentally different from an improvement heuristic (like 2-opt), which requires a complete solution as input for iterative refinement.
  • Comparison with the Appending-based Paradigm: The starting point of our work is precisely to overcome the limitations of the popular "appending-based paradigm" in existing NCO methods (as mentioned in the Introduction and Figure 1). "Appending" is a construction method, and so is "insertion." However, the latter offers greater flexibility and a larger solution space, which is the core of our exploration.
2. The Application of the Method: Distinguishing the Core Construction Model from the Inference Improvement Strategy
We would like to clarify that our adoption of the Insertion-based Local Reconstruction strategy during the inference stage (Section 3.3) is an advanced search strategy that uses our learned "construction model" as a core operator, rather than being the model itself.
This can be understood as a two-tiered structure:
  • Lower Level (Core Contribution): A learned, efficient construction heuristic model (L2C-Insert) that knows how to intelligently "insert" nodes.
  • Upper Level (Application Strategy): An improvement framework (akin to Large Neighborhood Search, LNS) that utilizes the lower-level construction model to iteratively "destroy and repair" a solution during inference to enhance its quality.
This practice of "embedding a learned construction model into an improvement framework" is one of the most recognized, state-of-the-art, and effective practices in the current NCO field (as referenced in the paper, e.g., LEHD [22], GLOP [40]). Our storyline does not confuse the two. Instead, it is:
  • We propose a new, more flexible construction model (L2C-Insert).
  • To validate its powerful performance, we follow the best practices in the field and use it as a potent "local reconstruction" tool during the inference stage.
3. Clear Technical Contributions
Based on the two points above, our technical contributions can be summarized more clearly as follows:
  • Paradigm Innovation: In the NCO field, we are the first to explore and successfully implement a learning-driven "insertion-based" construction framework, breaking the reliance of previous research on "appending-based" construction. This is an important supplement and a rethinking of the existing NCO construction methodology.
  • Novel Model Architecture and Training Mechanism: We designed a specific decoder structure (Section 3.1) to handle the complex "insertion" decision. It needs to simultaneously understand the intricate relationships between the node to be inserted, all unvisited nodes, and all possible insertion positions in the current partial solution. This is much more complex than an "appending" model, which only needs to focus on the last node. The corresponding supervised learning training scheme (Section 3.2) is also newly designed to fit the "insertion" task.
  • Synergistic Inference Strategy: We propose a distance-based destruction strategy (Section 3.3) that is highly synergistic with the "insertion" paradigm. As described in the paper, this strategy can more effectively identify and break "bad edges" that are spatially close but sequentially distant, a flexibility that traditional "sequence-based destruction strategies" lack. The experiment (Table 5) also proves its effectiveness. This demonstrates the powerful synergy between our proposed construction paradigm and inference strategy.
In summary, our work is not a simple application of a known "insertion heuristic" to NCO. Our core is a novel construction model, and its application during the inference stage is an advanced and reasonable improvement strategy. We believe that this clear hierarchical division and the technical innovations at each level together constitute the solid technical contribution of this paper.
Appendix J Analysis of Computational Complexity for Insertion-based and Appending-based heavy decoder-based NCO model
A formal analysis of computational complexity is essential for a thorough comparison. Here is a detailed comparison:
Computational Complexity of Appending-based NCO (L2C-Append)
The primary source of computational complexity of the Transformer-based model comes from the attention mechanism employed.
In a heavy decoder-based appending-based NCO model (e.g., LEHD, where heavy decoder means the decoder has six attention layers), the task at each construction step $t$ is to select one node from the ntn-t unvisited nodes. The decoder computes a context query and attends to the embeddings of all $n$ initial nodes to determine the next node to append. This attention mechanism has a complexity of $O((n-t)^{2}\cdot d)$, where $n$ is the total number of nodes and $d$ is the embedding dimension. This process is repeated for $n$ steps. Thus, the total complexity for constructing a single solution is $O(\frac{n^{3}}{3}\cdot d)$. In Big O notation, this is simplified to $O(n^{3}⋅d)$.
Computational Complexity of Insertion-based NCO (L2C-Insert)
Our proposed L2C-Insert framework also adopts a heavy decoder-based model. The task at each step $t$ is more complex. To make an informed decision, our decoder processes a much richer set of inputs simultaneously: the embedding of the current node to be inserted, the embeddings of all ntn−t remaining unvisited nodes, and the embeddings of all t1t−1 possible insertion positions. The total number of input embeddings processed by the self-attention layers at every step is therefore (1)+(nt)+(t1)=n(1)+(n-t)+(t-1)=n. The complexity of a single construction step is $O(n^{2}⋅d)$. This step is repeated $n$ times to build the full solution. Consequently, the total theoretical complexity is $O(n^{3}⋅d)$.
Discussion and Practical Implications
Theoretically, both heavy-decoder paradigms exhibit a cubic complexity of $O(n^{3}⋅d)$. However, the insertion-based paradigm is computationally more intensive as it consistently processes $n$ nodes per step, whereas the appending-based approach processes a decreasing number of nodes (nt)(n-t). This higher computational cost is a direct trade-off for the enhanced flexibility and expressive power of the insertion mechanism.
To ensure practical applicability to large-scale problems, we can employ a k-nearest neighbor (k-NN) strategy during inference. Instead of processing all $n$ contextual nodes, the decoder’s attention mechanism is constrained to only the k-nearest unvisited nodes and positions. This reduces the effective sequence length from $n$ to a much smaller, constant $k$ (e.g., $k=100$ in our paper), making the per-step complexity approximately $O(k^{2}⋅d)$.
Appendix K Clarification on the Definition of Target Position
We provide the following clarification to clarify the definition of the target position.
1. Why two visited nodes in the current partial solution determine the target position
During the process of the insertion construction, at each step, an unvisited node is inserted into a position of the partial solution constructed in the current step. Since 1) the nodes connected in the current partial solution are all visited nodes; 2) an insertion position is essentially an edge and is uniquely determined by the two nodes it connects, therefore, the only insertion position is determined by two visited points.
2. How the target position is determined
To make the process of identifying the target position clearer, we break it down into two steps. Given a labeled solution, a current partial solution, and a selected unvisited node to insert:
Step 1: Identify the circular sequence in the labeled solution
We first consider the complete labeled solution $\boldsymbol{\pi}^{*}$ and remove all unvisited nodes except for the currently selected one. This leaves us with a circular sequence containing only the selected unvisited node and the nodes already in the partial solution, preserving their original ordering from $\boldsymbol{\pi}^{*}$.
For example, suppose the labeled solution is $\boldsymbol{\pi}^{*}=(\pi_{1},\pi_{2},\pi_{3},\pi_{4},\pi_{5},\pi_{6},\pi_{7})$, the partial solution is $(\pi_{1},\pi_{3},\pi_{6},\pi_{7})$, and the selected unvisited node is $\pi_{2}$. Then the other unvisited nodes are $\{\pi_{4},\pi_{5}\}$. After removing $\{\pi_{4},\pi_{5}\}$ from $\boldsymbol{\pi}^{*}$, the resulting circular sequence is $(\pi_{1},\pi_{2},\pi_{3},\pi_{6},\pi_{7})$.
Step 2: Identify the target position in the current partial solution
Within this newly formed circular sequence, we can find the two nodes that are now adjacent to the selected unvisited node. These two nodes, which are part of the partial solution, define the target position for the insertion.
Continue the above example. In the circular sequence $(\pi_{1},\pi_{2},\pi_{3},\pi_{6},\pi_{7})$, the nodes adjacent to $\pi_{2}$ are $\{\pi_{1},\pi_{3}\}$. Since $\pi_{1}$ and $\pi_{3}$ are connected in the current partial solution, these two visited nodes determine the target position.
Appendix L Rationale and Principles of the Logarithmic Loss Function in Equation 3
The use of the negative log-likelihood loss function in Equation 3 is a standard and principled choice for our supervised learning framework, primarily because it is highly effective for gradient-based optimization.
Our model treats the selection of an insertion position at each construction step as a classification problem. The model’s decoder, via a softmax function, outputs a probability distribution $p_{\theta}$ over all valid insertion positions. Our objective is to train the model to assign the highest possible probability to the correct target position $\vec{e}^{\,*}$.
The logarithmic nature of this loss function $\mathcal{L}(\boldsymbol{\theta})=-\log p_{\boldsymbol{\theta}}\left(\vec{e}^{\,*}\mid\vec{e}_{\pi_{1:l}},\boldsymbol{\pi}^{*}\right)$ provides two key advantages for training:
Strong Error Correction: When the model assigns a very low probability to the correct target position (i.e., $p_{\theta}(\vec{e}^{\,*})\rightarrow$ 0 ), the loss $log\ p_{\theta}(\vec{e}^{\,*})$ approaches infinity. This results in a very large gradient, providing a strong signal to the optimizer to correct the model’s parameters significantly.
Stable Convergence: Conversely, as the model’s prediction for the correct class approaches 1, the loss approaches 0. This leads to smaller gradients, promoting stable convergence once the model learns to make correct predictions. This behavior effectively focuses the training effort on the most difficult examples.
Appendix M Visualization